Helen earns $7.50 an hour at her job, but if she works more than 35 hours in a week she is paid 1 1/2 times her regular salary for the overtime hours worked. One week her gross pay was $352.50. How many overtime hours did she work that week?
Found 2 solutions by stanbon, cattycassie77:
Answer by stanbon(48554)

Helen earns $7.50 an hour at her job, but if she works more than 35 hours in a week she is paid 1 1/2 times her regular salary for the overtime hours worked. One week her gross pay was $352.50. How many overtime hours did she work that week?
---------
Let "x" be the number of overtime hours she worked.
EQUATION:
Pay = regular pay + overtime pay
Pay= 7.50*35+(3/2)(7.50)x=352.50
62.5+11.25x=352.50
11.25x=341.25
x=30 1/3 hours.
Cheers,
Stan H.

Answer by cattycassie77(2)

(352.50-(35(7.40)))/(1.5x7.5)=X
X=8 hours overtime